A random sample of 9 fields of corn has a mean yield of 38.2 bushels per acre and standard deviation of 9.66 bushels per acre. Determine the 80%

confidence interval for the true mean yield. Assume the population is approximately normal.

Step 1 of 2 :

Find the critical value that should be used in constructing the confidence interval. Round your answer to three decimal places.
Answer #1

Solution :

Given that,

sample size = n = 9

Degrees of freedom = df = n - 1 = 9 - 1 = 8

t\alpha /2,df = 1.397

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 1.397 * (9.66 / \sqrt 9)

= 4.498

The 80% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

38.2 - 4.498 < \mu < 38.2 + 4.498

33.702 < \mu < 42.698

The 80%confidence interval for the true mean: (33.702 , 42.698)